A flat bar 75 mm wide by 9 mm thick and 1.2 m long is subjected to end moments of 202.5 N.m. Compute the deflection of the beam at the midspan if E = 200 000 MPa.
1
Expert's answer
2021-07-12T03:37:10-0400

Moment of inertia

"I= \frac{75*10^{-3}*(9*10^{-3})^3}{12}=4.5562*10^{-3}m^4\\
\delta = \frac{Ml^2}{8EI}\\
\delta = \frac{202.5*1.2^2}{8*200000*10^6*4.562*10^{-5}}\\
\delta =40*10^{-3}"